OpenStudy (anonymous):

It's been a long while since I've done some exercises on vectors, and I could use a refreshener on "how to". In this case I got the following x'= matrix . x. Then two vectors are given and I have to verify whether they are solutions for the system of DEs (expressed as a matrix). I'll post the matrix and vectors as a comment after this question post, but I prefer help on "how to" rather than a calculated answer.

OpenStudy (anonymous):

System of DEs: \[x'=\left[\begin{matrix}4 & -3 \\ 6 & -7\end{matrix}\right]x\] Verify whether the vectors \[x _{1}=\left(\begin{matrix}3e ^{2t} \\ 2e ^{2t}\end{matrix}\right)\] and \[x _{2}=\left(\begin{matrix}e ^{-5t} \\ 3e ^{-5t}\end{matrix}\right)\] are solutions for the system.

OpenStudy (anonymous):

Take x1. Calculate x1'.

OpenStudy (anonymous):

Plug x1 and x1' into the system. You get an identity?

OpenStudy (anonymous):

So when I calculate x'1 I get \[x _{1}^{'}=\left(\begin{matrix}6e ^{2t} \\ 4e ^{2t}\end{matrix}\right)\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Not sure how to write the matrix as DEs. They are not given except as the matrix.

OpenStudy (anonymous):

Our system is of the form \[X' = AX\]

OpenStudy (anonymous):

oh yeah... lol

OpenStudy (anonymous):

Identity isn't that a diagonal matrix with only 1s on the diagonal and the rest zero?

OpenStudy (anonymous):

We can think of it like this too\[\left(\begin{matrix}x_{11}' \\ x_{12}'\end{matrix}\right) = \left[\begin{matrix}4 & -3 \\ 6 & -7\end{matrix}\right]\left(\begin{matrix}x_{11} \\ x_{12}\end{matrix}\right)\]

OpenStudy (anonymous):

By identity, I mean something like x = x.

OpenStudy (anonymous):

When you calculate AX, you will get a vector.

OpenStudy (anonymous):

yeah, x'

OpenStudy (anonymous):

Does this vector equal X' ?

OpenStudy (anonymous):

But I just calculated x'1

OpenStudy (anonymous):

So x_1 is a solution then. Is x_2 a solution?

OpenStudy (anonymous):

If I can calculate it, it's a solution then?

OpenStudy (anonymous):

No. Sorry, I must have misunderstood. You calculated x_1', so now calculate Ax

OpenStudy (anonymous):

Where A is the coefficient matrix you were given.

OpenStudy (anonymous):

And without using x1 for x?

OpenStudy (anonymous):

With using x1.

OpenStudy (anonymous):

We are checking if x1 and x1' satisfies the equation, much like you would with a single DE

OpenStudy (anonymous):

Hmmm, I'm missing something here... I calculated x'1 by multiplying A with x1. And now I have to calculate Ax using x1. Isn't that doing the same thing twice?

OpenStudy (anonymous):

Lol.

OpenStudy (anonymous):

You have just shown that x1 is a solution.

OpenStudy (anonymous):

By calculate x1'

OpenStudy (anonymous):

I meant take x1, and take its derivative.

OpenStudy (anonymous):

So we would take the derivative of 3e^(2t) and 2e^(2t)

OpenStudy (anonymous):

so I derive 3e^2t and 2e^2t

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

And if that equals Ax, then x is a solution.

OpenStudy (anonymous):

I can see that on the spot, yeah, the derivative of x1 is x'1

OpenStudy (anonymous):

Yeah, just take [x1]d/dx, plug in x1 for x, and you see that they are the same immediately.

OpenStudy (anonymous):

yup d/dt of (3e^2t) = 6e^2t and d/dt of (2e^2t)=4e^2t

OpenStudy (anonymous):

So is x_2 a solution?

OpenStudy (anonymous):

So, by using both methods I prove I have the same result, and thus it's a solution.

OpenStudy (anonymous):

I'll check

OpenStudy (anonymous):

It's a solution because you derive an identity.

OpenStudy (anonymous):

The second vector isn't a solution I think x'2 calculated by multiplying the matrix with x2 = \[\left(\begin{matrix}e ^{-5t} \\ -3e ^{-5t}\end{matrix}\right)\] but the derivative of x2= \[\left(\begin{matrix}-5e^{-5t} \\ -15e ^{-5t}\end{matrix}\right)\]

OpenStudy (anonymous):

I think you should get <e^-5t, -15e^-5t> after the product, but nonetheless, it isn't a solution.

OpenStudy (anonymous):

Yeah, right, I multiplied the wrong elements for the second row

OpenStudy (anonymous):

Ok, thanks for the help... I have to test other stuff in this exercise but I'll try to see whether I can find it on my own first. Next step is using the Wronskian to prove the solutions are lineair independent.

OpenStudy (anonymous):

Fun stuff. Good luck.

OpenStudy (anonymous):

Hehehe. After systems of DEs I can start on multivariate analyses. I do find it fun stuff though... especially when you figure things out :-)

OpenStudy (anonymous):

BTW I just checked: x2 is a solution. I made the same calculation mistake for the first row as well. As I said, a bit rusty on the matrixes.

OpenStudy (anonymous):

Yes, you are right. 4 - 9. I didn't catch it at first either.

OpenStudy (anonymous):

A good exercise would be to find the eigenvalues of A and compute the solutions.

OpenStudy (anonymous):

Yeah, that's what I thought I was supposed to do at first.